OpenStudy (anonymous):

HELP!!!! Find the foci of 16y^2-36x^2-576=0. Then graph the hyperbole.

OpenStudy (loser66):

move -576 to the right side, what do you get?

OpenStudy (anonymous):

Positive 576

OpenStudy (loser66):

yes, now, divide both sides by 576, what do you get?

OpenStudy (anonymous):

1/36y^2-1/16x^2

OpenStudy (loser66):

=??

OpenStudy (anonymous):

=1?

OpenStudy (loser66):

\[\dfrac{y^2}{6^2}-\dfrac{x^2}{4^2}=1\] right?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

Hence, its b =6, a =4 ok?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

foci c, \(c=\sqrt{a^2+b^2}=??\)

OpenStudy (anonymous):

thanks can you stay with me for just a second i need to write this down

OpenStudy (anonymous):

ok is it square root of 52

OpenStudy (loser66):

You need know where c is. Your equation has y first, hence its graph is vertical, right? Hence foci will be (0, -c) and (0,c) . Yes for c =sqrt 52

OpenStudy (loser66):